The Base of the triangle is 4 inches longer than the length of one of the other sides. Use a system of equations to find the length of each side of the triangle. Perimeter =46 inches
Answer by ankor@dixie-net.com(22740)

The Base of the triangle is 4 inches longer than the length of one of the other sides. Use a system of equations to find the length of each side of the triangle. Perimeter = 46 inches
:
It doesn't say it is a right triangle, so there is no unique solution
:
designate the 3 sides as a, b, c
:
Write an equation for each statement
:
" Base of the triangle is 4 inches longer than the length of one of the other sides."
b = (a + 4)
and
a = (b - 4)
:
"Perimeter = 46 inches"
a + b + c = 46
Substitute (a+4) for b
a + (a+4) + c = 46
2a + c = 46 - 4
2a + c = 42
c = (42 - 2a)